May you help me with this problem:
3b+2c=46
5c+b=11
The instructions are to use linear combinations to solve the linear system. Then check your solution. Thank you for time.

Answer by stanbon(75887)

3b+2c=46
5c+b=11
----
Rearrange:
3b + 2c = 46
b + 5c = 11
---
Multiply 2nd equation by 3 to get:
3b + 2c = 46
3b + 15c = 33
------------------
Subtract top from bottom and solve for "c":
13c = -13
c = -1
----
Solve for "b":
5c+b = 11
5(-1) + b = 11
b = 11+5
b = 16
====================
cheers,
stan H.
============